# Data Processing for the OpenGPT-X Model Family

### v2

@@ 8.3.2 List of Curated Datasets
pile_pmc_extracts [32] EN JSONL Various [^36] This Pile part is extracted from PubMed where the author can decide between a variety of licenses. Medical 2.808.849 12.113 1%
ppc [75] PL XML CC BY 4.0 Law & Admin. 38.304 472.193 1%
principle_ga TXT GA, EN CC BY 4.0 Law & Admin. 11 25.776.098 3%
projekt_gutenberg EU24 TXT Custom Licence Books 60.912 3.374 1%
proof_pile EN JSONL Apache-2.0 Math 2.128.218 4.495 0%
rp_arxiv [15] EN JSONL Apache-2.0 Academic 1.503.469 10.202 3%
setimes_corpus [88] BG, BS, EL, EN, HR, MK, RO, SQ TXT CC BY-SA 3.0 News 246.206 69.279 7%
seimas [72] EN, LT XML / TMX CC BY 4.0 Law & Admin. 3540 47.319 0%
sk_court_decisions SVK JSONL CC0 1.0 Law & Admin. 1.647.736 2.174 74% [^37] Noisy documenta and conversion errors.
slwac [67] SL CONLLU CC BY-SA 4.0 Web 1.484.546 953.803 16% [^38] Low language score due to similar languages.
spanish_legal_corpora [40] ES TXT CC BY 4.0 Law & Admin. 15 1.383.749 0%
rp_stackexchange EN JSONL CC BY-SA 4.0 Forum 23.909.364 7.311 19% [^39] Short documents.
@@
starcoder_data [65] XX JSONL Various [^40] The Stack is a collection of source code from repositories with various licenses. Any use of all or part of the code gathered in The Stack must abide by the terms of the original licenses, including attribution clauses when relevant. Source Code 206.634.734 73.064 0%
swiss_judgment_prediction [74] DE, FR JSONL CC BY-SA 4.0 Law & Admin. 262.789 144.599 0%
swiss_legislation_corpus [46] DE, FR TXT CC BY-SA Law & Admin. 2 10.407 0%
swiss_policy_documents [36] DE, FR, IT PARQUET CC BY 4.0 Law & Admin. 417.923 561.574 0%
wacky [8] EN, DE, IT TXT CC BY-NC-SA 4.0 Web 8.571.072 6.440 0%
wikihow [56] EN CSV CC BY-NC-SA Knowledge Base 210.526 109.611 0%
wikibooks [30] EU24 XML CC BY-SA 4.0 Books 190.332 168.473 6%
wikinews [30] EU24 XML CC BY 2.5 News 197.265 46.936 3%
wikipedia [30] EU24 XML CC BY-SA 4.0 Knowledge Base 26.175.600 11.382 6%
wikiquote [30] EU24 XML CC BY-SA 4.0 Recreation 208.881 146.497 2%
wikisource [30] EU24 XML CC BY-SA 4.0 Books 428.157 798.322 7%
wikivoyage [30] EU24 XML CC BY-SA 4.0 Culture 71.905 66.278 4%
wikiversity [30] EU24 XML CC BY-SA 4.0 Books 61.397 65.482 0%

### v4

@@ 8.3.1 Filtering
pile_pmc_extracts [32] EN JSONL Various [^30] This Pile part is extracted from PubMed where the author can decide between a variety of licenses. Medical 2.808.849 12.113
ppc [75] PL XML CC BY 4.0 Law & Admin. 38.304 472.193
principle_ga TXT GA, EN CC BY 4.0 Law & Admin. 11 25.776.098
projekt_gutenberg EU24 TXT Custom Licence Books 60.912 3.374
proof_pile EN JSONL Apache-2.0 Math 2.128.218 4.495
rp_arxiv [15] EN JSONL Apache-2.0 Academic 1.503.469 10.202
setimes_corpus [88] BG, BS, EL, EN, HR, MK, RO, SQ TXT CC BY-SA 3.0 News 246.206 69.279
seimas [72] EN, LT XML / TMX CC BY 4.0 Law & Admin. 3540 47.319
sk_court_decisions SVK JSONL CC0 1.0 Law & Admin. 1.647.736 2.174
slwac [67] SL CONLLU CC BY-SA 4.0 Web 1.484.546 953.803
spanish_legal_corpora [40] ES TXT CC BY 4.0 Law & Admin. 15 1.383.749
rp_stackexchange EN JSONL CC BY-SA 4.0 Forum 23.909.364 7.311
@@
starcoder_data [65] $X$ JSONL Various [^31] The Stack is a collection of source code from repositories with various licenses. Any use of all or part of the code gathered in The Stack must abide by the terms of the original licenses, including attribution clauses when relevant. Source Code 206.634.734 73.064
swiss_judgment_prediction [74] DE, FR JSONL CC BY-SA 4.0 Law & Admin. 262.789 144.599
swiss_legislation_corpus [46] DE, FR TXT CC BY-SA Law & Admin. 2 10.407
swiss_policy_documents [36] DE, FR, IT PARQUET CC BY 4.0 Law & Admin. 417.923 561.574
wacky [8] EN, DE, IT TXT CC BY-NC-SA 4.0 Web 8.571.072 6.440
wikihow [56] EN CSV CC BY-NC-SA Knowledge Base 210.526 109.611
wikibooks [30] EU24 XML CC BY-SA 4.0 Books 190.332 168.473
wikinews [30] EU24 XML CC BY 2.5 News 197.265 46.936
wikipedia [30] EU24 XML CC BY-SA 4.0 Knowledge Base 26.175.600 11.382
wikiquote [30] EU24 XML CC BY-SA 4.0 Recreation 208.881 146.497
wikisource [30] EU24 XML CC BY-SA 4.0 Books 428.157 798.322
wikivoyage [30] EU24 XML CC BY-SA 4.0 Culture 71.905 66.278
wikiversity [30] EU24 XML CC BY-SA 4.0 Books 61.397 65.482